OpenStudy (anonymous):

Width of rectangle is 8 more the length. The perimeter is 56. Help me solve with steps? Find the dimentions Find the area?

OpenStudy (anonymous):

lets say the length is L then the width is L+8 the perimeter is 2(length+width) so, 56 = 2(L+L+8) solve for L. that is the length. width is 8 more than length. area is length times width

OpenStudy (blacksteel):

Alright, let's let the width be w and the length be l. We know that w = l + 8 and we know the the perimeter is 2w + 2l = 56 Then we can substitute the first equation into the second to get: 2(l + 8) + 2l = 56 2l + 16 + 2l = 56 4l = 40 l = 10. Then we can use the first equation again to get w = 10 + 8 = 18. So the dimensions are 10x18. The area is just the length times the width, so 18 * 18 or 180.

OpenStudy (blacksteel):

*10*18 or 180.

myininaya (myininaya):

W=8+L 2L+2W=56 ----------------- 2L+2(8+L)=56 4L+16=56 4L=40 L=10 so W=8+10=18 A=LW=18(10)=180

OpenStudy (anonymous):

the area would be 180 good thanks